### v1

@@ 5 The local lemma and graph colouring
A proper $\hat{L}$-colouring $f$ of $G$ corresponds to an independent set of size $|V_{G}|$ in $H$ – namely, the set $I=\{(u,f(u)):u\in V_{G}\}$. Because of the partition of $|V_{H}|$ into $|V_{G}|$ cliques given by $L$, the largest possible size of an independent set in $H$ is $|V_{G}|$, and so we have reduced the problem of that of showing that $G$ is $\hat{L}$-colourable to showing that $\alpha(H)=|V_{G}|$. For convenience, we define an independent set $I$ in the cover $H$ to be saturating if it saturates the bound $|I|\leq|V_{G}|$. (Such an $I$ is also referred to as an independent transversal.) One might hope that saturating independent sets can be found with local occupancy via Theorem 3.5, but this turns out to be a naïve hope in general.
@@
    $A$ is mutually independent of a set of all but at most $d$ of the other events.
@@
Proof of Theorem 5.1.
Let $S$ be formed by selecting, for each $u\in V_{G}$, a vertex $x\in L(u)$ uniformly at random. For an edge $xy$ in $H$, let $B_{xy}$ be the bad event that $xy\in H[S]$. Note that $S$ is a saturating independent set if and only if no bad event occurs. Also, if $x\in L(u)$ and $y\in L(v)$, then
@@
and $B_{xy}$ is independent of $B_{zw}$ unless at least one of $z$ and $w$ lies in $L(u)\cup L(v)$. Thus, in the dependency graph $B_{xy}$ has degree at most $d|L(u)|+d|L(v)|-1$ (excluding $B_{xy}$ itself). By the Lovász local lemma, the proof is complete if, for all $uv\in E_{G}$ and $xy\in E_{H}$ with $x\in L(u)$ and $y\in L(v)$, we have
@@
but this follows from the lower bound on the list sizes. ∎
@@
While Theorem 5.1 is a good start, it is not so powerful on its own. For example, we cannot straightforwardly recover the greedy bound, which states $\chi(G)\leq\Delta+1$. If we consider a cover for proper $(\Delta+1)$-colourings of the complete graph $K_{\Delta+1}$, then every colour-degree is $\Delta$ and there is no version of Theorem 5.1 for these parameters.
@@
For locally sparse $G$, this strategy has been useful since the 1990s: both Kim [52] (for $G$ being of girth $5$) and Johansson [46] (for $G$ being triangle-free) found an appropriate set $I$ by nibbling, an approach covered in Chapter 8. It was a breakthrough in 2019 when Molloy [56] showed that a suitable $I$ can be found by using entropy compression to analyze an algorithm for generating a random independent set. Bernshteyn [10] subsequently showed that Molloy’s argument can be cast in terms of analyzing a uniformly random independent set in the cover using the so-called ‘lopsided’ Lovász local lemma. Later work [29] involving the authors showed that the conditions under which Bernshteyn’s analysis applies follow from local occupancy.
@@
The idea of considering a random independent set $Y$ in $H$, distributed according to the hard-core model at fugacity $\lambda$, is the main conceptual breakthrough that connects local occupancy with proper colourings. Only for very large $\lambda$ might we expect to gain information about saturating independent sets in $H$, but for the best results we are forced to handle smaller values of $\lambda$ and find a smaller independent set which allows for an application of Theorem 5.1. This is the reason for some ‘mild additional technical conditions’ in the colouring aspects of local occupancy mentioned in the Introduction. We make the following standing assumptions for the remainder of this section:
@@
  • $X$ is a random independent set of $G$ chosen according to the hard-core model at fugacity $\lambda$, and $X$ has local $(\beta,\gamma)$-occupancy as per Definition 3.3;
@@
This section is focused on showing that, with suitable parameter choices, there is positive probability that the leftover – that is, the cover $\mathscr{H}_{Y}$ together with the graph $G_{Y}$ – satisfies the conditions of Theorem 5.1. Since we need only to know about $L_{Y}(u)$ and $\deg^{*}(x)$ for $x\in L(u)$ when $u\in V(G_{Y})$, we can afford to condition on $u\in V(G_{Y})$ in the probabilistic analysis. That is why we can fix $u$ and consider the random independent set $Y^{\prime}$, which by the spatial Markov property (Theorem 3.1) is distributed according to the hard-core model on $H^{\prime}=H-L(u)$.
@@
We note that $x\in L_{Y^{\prime}}(u)$ if and only if $Y^{\prime}\cap N_{H}(x)=\varnothing$, and also that the set $N_{H}(x)\cap Y^{\prime}$ of occupied neighbours of $x$ is simply the set of occupied vertices in $F$. Then the spatial Markov property (Theorem 3.1) gives
@@
as required for the equation (9.7). But $H^{\prime}[N_{H}(x)]$ is isomorphic to an induced subgraph of $G[N(u)]$, and hence $F$ is too, for any realization of $Y^{\prime}\setminus N_{H}(x)$. This permits an application of the local occupancy condition that we assume for $X$, and the inequality (9.8) now follows from the inequality (9.4). ∎
Theorem 5.2 clearly requires the additional power of the condition in Definition 3.3, in comparison with the weaker conditions in Theorem 3.4. Under the weaker conditions for $X$, it is unclear how one could conclude anything useful about the hard-core model on a cover of $G$. For the remainder of this section, we make no further reference to $X$, save for the consequences that it has for $Y^{\prime}$ as per Theorem 5.2.
@@
Due to the rather general setup, it is challenging to interpret the implications of Theorem 5.3. We can substitute in bounds on the $Z_{F}(\lambda)$ terms subject to local information. In the setting discussed in the Introduction, we needed to allow for complete graphs $F$ and hence we could only ensure that $Z_{F}(\lambda)\geq 1+d\lambda$ for $|V(F)|=d$. But in the setting of triangle-free graphs, $F$ contains no edges and we may use the stronger bound $Z_{F}(\lambda)\geq(1+\lambda)^{d}$. In either case, we see that Theorem 5.3 provides an understanding of the first moment of quantities relevant to Theorem 5.1.
Proof of Theorem 5.3.
The first inequality follows from Theorem 5.2. For, if we sum the inequality (9.8) over all vertices $x\in L(u)$ we get
@@
The same bounds also hold subject to the additional conditioning on $Y^{\prime}\setminus L(N(u))$ because all we need is a property of all subgraphs $F$ that arise as subgraphs of $G[N(u)]$. In particular, Theorem 5.2 holds subject to such conditioning. ∎
@@
(YNH(y) for all $\mathbb{P}\big{(}\text{$Y^{\prime}\cap N_{H}(y)\neq\varnothing$ for all $y\in S\mid Y^{\prime}\cap N_{H}(x)=\varnothing$}\big{)}\geq\mathbb{P}(S\cap L_{Y^{\prime}}(u)=\varnothing),$
which holds because the sets $N_{H}(x)\setminus L(u)$ and $N_{H}(S)\setminus L(u)$ are disjoint.
The second bound follows from the inequality (9.9) and a union bound over the $q$ elements $x\in L(u)$. ∎
@@
Proof.
We use the Lovász local lemma to show that there is an independent set $I$ in $H$ for which the conditions of Theorem 5.1 hold in $\mathscr{H}_{I}$. For $u\in V_{G}$, let $A_{u}$ be the bad event that $u\in V(G_{Y})$ and $|L_{Y}(u)|<2ed$, and let $B_{u}$ be the bad event that $u\in V(G_{Y})$ and there exists some $x\in L_{Y}(u)$ with $\deg^{*}_{\mathscr{H}_{Y}}(x)>d$. We consider the family of events $C_{u}=A_{u}\cup B_{u}$ in the local lemma. Note that $C_{u}$ is independent of $C_{v}$ for any $v\in V_{G}\setminus N^{3}[u]$. This is because the bad events are all properties of the list $L_{Y}(u)$ of $u$ in the leftover cover, and this is fixed by the intersection of $Y$ with $N_{H}(L(u))$. If the distance in $G$ between $u$ and $v$ is at least $4$, then $N_{H}(L(u))$ and $N_{H}(L(v))$ are disjoint. Since $|N^{3}[u]|\leq\Delta^{3}$, it suffices to show that $\mathbb{P}(A_{u}),\mathbb{P}(B_{u})\leq 1/(2e\Delta^{3})$.
The bound $\mathbb{P}(A_{u})\leq 1/(2e\Delta^{3})$ follows from Theorem 5.4, the condition (9.10) and the lower bound on $d$. The bound $\mathbb{P}(B_{u})\leq 1/(2e\Delta^{3})$ follows from Theorem 5.4 and the condition (9.11). ∎
The astute reader may already have noticed that the setup above straightforwardly generalizes to list colouring, and so we can obtain the following results mainly as consequences of the local occupancy analyses of Section 4. The first of these is Molloy’s theorem [56], while the second and third of these are smooth generalizations of Molloy’s theorem as shown in [29].
Corollary 5.6
@@
Proof.
We established the fractional variants of these upper bounds in Section 4. What remains is to show that the ‘mild’ additional conditions of Theorem 5.5 hold, and to study the asymptotic growth of the resulting bounds on the list chromatic number. In each case we have a bound $\rho$ on the average degree in any subgraph of a neighbourhood in $G$, and we can thus apply Theorems 4.4 and 4.5. For convenience, we have stated conditions which guarantee that $\rho=o(\Delta)$.
@@
The result now follows because we can take $\rho=\sqrt{2t}$ for part (b) and $\rho=k-3$ for part (c). The final detail is to argue that $\log(\Delta/\sqrt{2t})\sim\log(\Delta/\sqrt{1+t})$ and $\log(\Delta/(k-3))\sim\log(\Delta/k)$ under the present assumptions, and this is simple asymptotic analysis. ∎
@@ 6 The occupancy fraction of regular graphs
Proof.
Write $n=|V|$ and recall that $\alpha_{G}(\lambda)=\frac{1}{n}\sum_{u\in V}\mathbb{P}(u\in X)$, where $X$ is distributed according to the hard-core model on $G$ at fugacity $\lambda$. We reveal $Y_{u}=|X\cap N(u)|$ and observe that
@@
by a calculation from the Introduction. Here, the first equality follows from the spatial Markov property and the second equality uses the fact that each vertex appears $d$ times in the double sum $\sum_{u\in V}\sum_{v\in N(u)}\mathbb{P}(v\in X)=\sum_{u\in V}\mathbb{E}Y_{u}$. We can further simplify our expressions for $\alpha_{G}(\lambda)$, by defining $Y$ to be $Y_{u}$ for a uniformly random choice of $u$ and writing $q_{i}=\mathbb{P}(Y=i)$. By linearity of expectation, we have
@@
The main idea now is to relax the above optimization problem, which formally is over all probability vectors $\mathbf{q}=(q_{0},q_{1},\dotsc,q_{d})$ that arise as the distribution of $Y$ on a $d$-regular graph, to a linear program over all real vectors $\mathbf{q}\in\mathbb{R}^{d+1}$. For high-quality bounds, we need a definition of $Y$ that reveals enough graph structure to capture the class of interest, and we want to include as many valid linear constraints as possible in order to reduce the difference between the true optimum and the optimum of the relaxation. Thus far, we are missing some important constraints. Rather like the constraint used to establish the condition in (9.6), for $i\geq 1$ we can take any realization of the random independent set $X$ and random vertex $u$ which yields $Y=i$, and form $X^{\prime}$ by removing any vertex in $X\cap N(u)$. This gives a pair $(X^{\prime},u)$ for which we would find $Y=i-1$. The important thing here is to take into account both any overcounting and the change in size $|X|=|X^{\prime}|+1$ which entails a factor of $\lambda$ in the probabilities. Each such $X^{\prime}$ can be found by removing a vertex from at most $d-i-1$ different $X$s, and each $X$ can have a vertex removed in $i$ ways, yielding the constraint
@@
We start by proving that if a vector $\mathbf{q}$ achieves the maximum, then the inequality (9.12) must hold with equality for $2\leq i\leq d$; this follows by a simple re-weighting argument. If for $i\geq 2$ we have strict inequality, then for a small enough $\varepsilon>0$ we can increase $q_{0}$ by $\varepsilon$, decrease $q_{i-1}$ by $\left(\frac{d\lambda}{1+\lambda}+i\right)\varepsilon$ and increase $q_{i}$ by $\left(\frac{d\lambda}{1+\lambda}+i-1\right)\varepsilon$ while maintaining all the constraints. But since we have increased $q_{0}$ we have also increased the objective value, which is a contradiction. It is now a simple, if tedious, calculation to show that there is precisely one vector $\mathbf{q}$ that achieves equality in (9.12) for $2\leq i\leq d$, whilst satisfying all the other constraints, and that the vector $\mathbf{q}$ is exactly the distribution of $Y$ when the graph is $K_{d,d}$.
Recalling (9.3), we then integrate $\mathbb{E}|X|/\lambda$ to obtain an estimate on $\log Z_{G}(\lambda)$ for the second assertion. ∎
The above ideas can be applied to a setting with much more challenging graph structure. Recall that, given a graph $G=(V,E)$, the line graph $L(G)$ is a graph with vertex-set $E$, where we connect two edges of $G$ in $L(G)$ if they share a common vertex. As the line graph of a $d$-regular simple graph is $2(d-1)$-regular, it turns out that if we tolerate a slight change in the parameter we can then ask for a version of Theorem 6.1 over the class of line graphs of regular graphs. For $d\geq 3$, the graph $K_{d,d}$ is not a line graph, and so we might hope for an improved upper bound. The following tight result from [26] establishes that $L(K_{d,d})$ is the extremal graph.

### v2

@@ 5 The local lemma and graph colouring
A proper $\hat{L}-$colouring $f$ of $G$ corresponds to an independent set of size $|V_{G}|$ in $H$ – namely, the set $I=\{(u,f(u)):u\in V_{G}\}$. Because of the partition of $|V_{H}|$ into $|V_{G}|$ cliques given by $L$, the largest possible size of an independent set in $H$ is $|V_{G}|$, and so we have reduced the problem of that of showing that $G$ is $\hat{L}-$colourable to showing that $\alpha(H)=|V_{G}|$. For convenience, we define an independent set $I$ in the cover $H$ to be saturating if it saturates the bound $|I|\leq|V_{G}|$. (Such a set $I$ is also referred to as an independent transversal.) One might hope that saturating independent sets can be found with local occupancy via Theorem 3.3, but this turns out to be a naïve hope in general.
@@
    AA is mutually independent of a set of all but at most $d$ of the other events.
@@
To see how Theorem 5.1 holds, let $S$ be formed by selecting, for each $u\in V_{G}$, a vertex $x\in L(u)$ uniformly at random. For an edge $xy$ in $H$, let $B_{xy}$ be the bad event that $xy\in H[S]$. Note that $S$ is a saturating independent set if and only if no bad event occurs. Also, if $x\in L(u)$ and $y\in L(v)$, then
@@
and $B_{xy}$ is independent of $B_{zw}$ unless at least one of $z$ and $w$ lies in $L(u)\cup L(v)$. Thus, in the dependency graph, $B_{xy}$ has degree at most d|L(u)|+d|L(v)|1d|L(u)|+d|L(v)|-1 (excluding $B_{xy}$ itself). By the Lovász local lemma, the proof is complete if, for all $uv\in E_{G}$ and $xy\in E_{H}$ with $x\in L(u)$ and $y\in L(v)$, we have
@@
but this follows from the lower bound on the list sizes, establishing the result.
@@
While Theorem 5.1 is a good start, it is not so powerful on its own. For example, we cannot straightforwardly recover the greedy bound, which states that $\chi(G)\leq\Delta+1$. If we consider a cover for proper $(\Delta+1)$-colourings of the complete graph $K_{\Delta+1}$, then every colour-degree is $\Delta$ and there is no version of Theorem 5.1 for these parameters.
@@
For a locally sparse graph $G$, this strategy has been useful since the 1990s: both Kim [57] (for $G$ of girth $5$) and Johansson [51] (for $G$ triangle-free) found an appropriate set $I$ by nibbling, an approach covered in Chapter 7. It was a breakthrough in 2019 when Molloy [62] showed that a suitable set $I$ can be found by using entropy compression to analyze an algorithm for generating a random independent set. Bernshteyn [12] subsequently showed that Molloy’s argument can be cast in terms of analyzing a uniformly random independent set in the cover using the so-called ‘lopsided’ Lovász local lemma. Later work [33] involving the authors showed that the conditions under which Bernshteyn’s analysis applies follow from local occupancy.
@@
The idea of considering a random independent set $Y$ in $H$, distributed according to the hard-core model at fugacity $\lambda$, is the main conceptual breakthrough that connects local occupancy with proper colourings. Only for very large $\lambda$ might we expect to gain information about saturating independent sets in $H$, but for the best results we are forced to handle smaller values of $\lambda$ and find a smaller independent set which allows for an application of Theorem 5.1. This is the reason for some ‘mild additional technical conditions’ in the colouring aspects of local occupancy mentioned in Section 1. We make the following standing assumptions for the remainder of this section:
@@
  • $X$ is a random independent set of $G$ chosen according to the hard-core model at fugacity $\lambda$, and $X$ has local $(\beta,\gamma)$-occupancy as per the Key Definition;
@@
This section is focused on showing that, with suitable parameter choices, there is positive probability that the leftover – that is, the cover $\mathscr{H}_{Y}$ together with the graph $G_{Y}$ – satisfies the conditions of Theorem 5.1. Since we only need to know about $L_{Y}(u)$ and $\deg^{*}(x)$ for $x\in L(u)$ when $u\in V(G_{Y})$, we can afford to condition on $u\in V(G_{Y})$ in the probabilistic analysis. That is why we can fix $u$ and consider the random independent set $Y^{\prime}$, which by the spatial Markov property (Theorem 3.1) is distributed according to the hard-core model on $H^{\prime}=H-L(u)$.
@@
To see how Theorem 5.2 holds, we note that $x\in L_{Y^{\prime}}(u)$ if and only if $Y^{\prime}\cap N_{H}(x)=\varnothing$, and also that the set $N_{H}(x)\cap Y^{\prime}$ of occupied neighbours of $x$ is simply the set of occupied vertices in $F$. Then the spatial Markov property (Theorem 3.1) gives
@@
as required for the equation (8.8). But $H^{\prime}[N_{H}(x)]$ is isomorphic to an induced subgraph of $G[N(u)]$, and hence $F$ is too, for any realization of $Y^{\prime}\setminus N_{H}(x)$. This permits an application of the local occupancy condition that we assume for $X$, and the inequality (8.9) now follows from the inequality (8.5).
Theorem 5.2 clearly requires the additional power of the condition in the Key Definition, in comparison with the weaker conditions in Theorem 3.2. Under the weaker conditions for $X$, it is unclear how one could conclude anything useful about the hard-core model on a cover of $G$. For the remainder of this section, we make no further reference to $X$, save for the consequences that it has for $Y^{\prime}$ as per Theorem 5.2.
@@
Due to the rather general setup, it is challenging to interpret the implications of Theorem 5.3. We can substitute in bounds on the $Z_{F}(\lambda)$ terms subject to local information. In the setting discussed in Section 1, we needed to allow for complete graphs $F$ and hence we could only ensure that $Z_{F}(\lambda)\geq 1+d\lambda$ for $|V(F)|=d$. But in the setting of triangle-free graphs, $F$ contains no edges and we may use the stronger bound $Z_{F}(\lambda)\geq(1+\lambda)^{d}$. In either case, we see that Theorem 5.3 provides an understanding of the first moment of quantities relevant to Theorem 5.1.
To see why Theorem 5.3 holds, the first inequality follows from Theorem 5.2. For, if we sum the inequality (8.9) over all vertices $x\in L(u)$ we get
@@
The same bounds also hold subject to the additional conditioning on $Y^{\prime}\setminus L(N(u))$ because all we need is a property of all subgraphs $F$ that arise as subgraphs of $G[N(u)]$. In particular, Theorem 5.2 holds subject to such conditioning, and this establishes Theorem 5.3.
@@
$\mathbb{P}\big{(}\text{$Y^{\prime}\cap N_{H}(y)\neq\varnothing$ for all $y\in S\mid Y^{\prime}\cap N_{H}(x)=\varnothing$}\big{)}\geq\mathbb{P}(S\cap L_{Y^{\prime}}(u)=\varnothing).$
In the triangle-free case this holds because the sets $N_{H}(x)\setminus L(u)$ and $N_{H}(S)\setminus L(u)$ are disjoint, and in the general case negative correlation can be established with a more involved argument.
The second bound in Theorem 5.4 follows from the inequality (8.10) and a union bound over the $q$ elements $x\in L(u)$.
@@
To see this, we use the Lovász local lemma to guarantee an independent set $I$ in $H$ for which the conditions of Theorem 5.1 hold in $\mathscr{H}_{I}$. For $u\in V_{G}$, let $A_{u}$ be the bad event that $u\in V(G_{Y})$ and $|L_{Y}(u)|<2ed$, and let $B_{u}$ be the bad event that $u\in V(G_{Y})$ and there exists some $x\in L_{Y}(u)$ with $\deg^{*}_{\mathscr{H}_{Y}}(x)>d$. We consider the family of events $C_{u}=A_{u}\cup B_{u}$ in the local lemma. Note that $C_{u}$ is independent of $C_{v}$ for any $v\in V_{G}\setminus N^{3}[u]$. This is because the bad events are all properties of the list $L_{Y}(u)$ of $u$ in the leftover cover, and this is fixed by the intersection of $Y$ with $N_{H}(L(u))$. If the distance in $G$ between $u$ and $v$ is at least $4$, then $N_{H}(L(u))$ and $N_{H}(L(v))$ are disjoint. Since $|N^{3}[u]|\leq\Delta^{3}$, it suffices to show that $\mathbb{P}(A_{u}),\mathbb{P}(B_{u})\leq 1/(2e\Delta^{3})$.
The bound $\mathbb{P}(A_{u})\leq 1/(2e\Delta^{3})$ follows from Theorem 5.4, the condition (8.11) and the lower bound on $d$. The bound $\mathbb{P}(B_{u})\leq 1/(2e\Delta^{3})$ follows from Theorem 5.4 and the condition (8.12). This establishes Theorem 5.5.
The astute reader may already have noticed that the setup above straightforwardly generalizes to list colouring, and so we can obtain the following results mainly as consequences of the local occupancy analyses of Section 4. The first of these is Molloy’s theorem [62], while the second and third are smooth generalizations of Molloy’s theorem as shown in [33].
Corollary
@@
We established the fractional variants of these upper bounds in Section 4. To see the above corollary, what remains is to show that the ‘mild’ additional conditions of Theorem 5.5 hold, and to study the asymptotic growth of the resulting bounds on the list chromatic number. In each case we have a bound $\rho$ on the average degree in any subgraph of a neighbourhood in $G$, and we can thus apply Theorems 4.3 and 4.4. For convenience, we have stated conditions which guarantee that $\rho=o(\Delta)$.
@@
The result now follows because we can take $\rho=\sqrt{2t}$ for part (b) and $\rho=k-3$ for part (c). The final detail is to argue that $\log(\Delta/\sqrt{2t})\sim\log(\Delta/\sqrt{1+t})$ and $\log(\Delta/(k-3))\sim\log(\Delta/k)$ under the present assumptions, and this is simple asymptotic analysis.
@@ 6 The occupancy fraction of regular graphs
To see why this result holds, we write $n=|V|$ and recall that $\alpha_{G}(\lambda)=\frac{1}{n}\sum_{u\in V}\mathbb{P}(u\in X)$, where $X$ is distributed according to the hard-core model on $G$ at fugacity $\lambda$. We reveal $Y_{u}=|X\cap N(u)|$ and observe that
@@
by a calculation from Section 1. Here, the first equality follows from the spatial Markov property and the second equality uses the fact that each vertex appears $d$ times in the double sum $\sum_{u\in V}\sum_{v\in N(u)}\mathbb{P}(v\in X)=\sum_{u\in V}\mathbb{E}Y_{u}$. We can further simplify our expressions for $\alpha_{G}(\lambda)$, by defining $Y$ to be $Y_{u}$ for a uniformly random choice of $u$ and writing $q_{i}=\mathbb{P}(Y=i)$. By linearity of expectation, we have
@@
The main idea now is to relax the above optimization problem, which formally is over all probability vectors $\mathbf{q}=(q_{0},q_{1},\dotsc,q_{d})$ that arise as the distribution of $Y$ on a $d$-regular graph, to a linear program over all real vectors $\mathbf{q}\in\mathbb{R}^{d+1}$. For high-quality bounds, we need a definition of $Y$ that reveals enough graph structure to capture the class of interest, and we want to include as many valid linear constraints as possible in order to reduce the difference between the true optimum and the optimum of the relaxation. Thus far, we are missing some important constraints. Rather like the constraint used to establish the condition in (8.7), for $i\geq 1$ we can take any realization of the random independent set $X$ and random vertex $u$ which yields $Y=i$, and form $X^{\prime}$ by removing any vertex in $X\cap N(u)$. This gives a pair $(X^{\prime},u)$ for which we would find Y=i1Y=i-1. The important thing here is to take into account both any overcounting and the change in size $|X|=|X^{\prime}|+1$ which entails a factor of $\lambda$ in the probabilities. Each such $X^{\prime}$ can be found by removing a vertex from at most di1d-i-1 different $X$s, and each $X$ can have a vertex removed in $i$ ways, yielding the constraint
@@
We start by proving that if a vector $\mathbf{q}$ achieves the maximum, then the inequality (8.13) must hold with equality for $2\leq i\leq d$; this follows by a simple re-weighting argument. If for $i\geq 2$ we have strict inequality, then for a small enough $\varepsilon>0$ we can increase $q_{0}$ by $\varepsilon$, decrease $q_{i-1}$ by $\left(\frac{d\lambda}{1+\lambda}+i\right)\varepsilon$ and increase $q_{i}$ by $\left(\frac{d\lambda}{1+\lambda}+i-1\right)\varepsilon$, while maintaining all the constraints. But since we have increased $q_{0}$ we have also increased the objective value, which is a contradiction. It is now a simple, if tedious, calculation to show that there is precisely one vector $\mathbf{q}$ that achieves equality in (8.13) for $2\leq i\leq d$, while satisfying all the other constraints, and that the vector $\mathbf{q}$ is exactly the distribution of $Y$ when the graph is $K_{d,d}$.
Recalling (8.3), we then integrate $\mathbb{E}|X|/\lambda$ to obtain an estimate on $\log Z_{G}(\lambda)$ for the second assertion. This establishes Theorem 6.1.
The above ideas can be applied to a setting with much more challenging graph structure. Recall that, given a graph $G=(V,E)$, the line graph $L(G)$ is a graph with vertex-set $E$, where we connect two edges of $G$ in $L(G)$ if they share a common vertex. As the line graph of a $d$-regular simple graph is 2(d1)2(d-1)-regular, it turns out that if we tolerate a slight change in the parameter we can then ask for a version of Theorem 6.1 over the class of line graphs of regular graphs. For $d\geq 3$, the graph $K_{d,d}$ is not a line graph, and so we might hope for an improved upper bound. The following tight result from [30] establishes that $L(K_{d,d})$ is the extremal graph.